Week of 170313

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Kate (WLCG, chair), Julia A (WLCG), Vincent (security), Gavin (computing), Andrea M (middleware), Belinda (storage), Mark Slater (LHCb), Maarten (WLCG, Alice)
  • remote: Luca Lama, Elena Corni (CNAF), David Bouvet (IN2P4), FaHui Lin (ASGC), John K (RAL), Dave Mason (FNAL), Sang-Un (KISTI), Tom W (NL-T1), Di Qing (TRIUMF), Victor Z. (JINR), Pepe M. (PIC), Dimitri (KIT)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Software and Computing Week, apologies for not connecting -PeterLove
    • Production at high levels (300-350k cores)
    • CERN-wide power cut affected Sim@P1 (HLT farm)
    • bad job monitoring in dashboard, ticket raised DASHB-2991
    • MC15 evgen jobs new memory monitor shows high disk IO at 20MB/s read and write, affecting some sites
    • Reprocessing started Friday and Frontier/squid issues over the weekend, ATLAS looking at software details and also RAL ticket ongoing GGUS-127079
    • ATLAS Robot certificates suspended from VOMS, various emails have been sent now fixed, need to followup re-occurring VOMS server issues

  • CMS reports ( raw view) -
    • I could be travelling @ the meeting times, please forgive my absence (if it will happen!)
    • CMS General:
      • production + analysis ongoing; sites full in the second part of the week, a bit less in the first part
    • No real issues, just a Kibana "blackout" on Sat/Sun. Being understood whether it is a kibana issue / CMSWEB simply not sending monitoring data
    • Nothing else crucial to report.

  • ALICE -
    • CERN: condorce01 lost 8k ALICE jobs on Sat (GGUS:127076)

  • LHCb reports ( raw view) -
    • Activity
      • MC Simulation, Stripping campaign now started so tape systems will start to be hit
    • Site Issues
      • T0:
        • ALARM ticket GGUS:126874 about users running out of AUP signature validity. User AUP validity overwritten with admin rights. No updates since 2nd March - any more news on fixes? Maarten commented that there is a number of issues to be fixed. It would be good to remove the ALARM.
        • Ready for network outages on Wednesday morning - Thanks for shifting the DT from 22nd to 15th as well!
      • T1:
        • FZK: Some ongoing issues with submission timeouts to the new ARC CEs with arc-2-kit not working at all (GGUS:127075).
Maarten: Issue is also observed by Alice but doesn't affect operations.

Sites / Services round table:

  • ASGC: ntr
  • BNL: dCache upgrade (from 2.10 to 3.0) is scheduled for 03/22 and 03/23. BNL computing farm will be closed to new jobs during that time period. Announcement is posted to OSG OIM and ATLAS ELOG.
  • CNAF: ntr
  • EGI: nc
  • FNAL:ntr
  • IN2P3:
    • Router for LHCOPN has been changed on March 8th.
    • Scheduled maintenance on next Tuesday March 21st.
      • 06:00 - 07:00 UTC : site will be disconnected as external connection router will be changed.
      • whole day : CEs, SEs, Frontier servers and XrootD redirectors will be down
  • JINR: LAN storm 11-03-2017 due to lost power in one building and border gateway switch misconfiguration. Fixed.
  • KISTI: ntr
  • KIT: ntr (ARC CE issues from time to time, developers are checking, bug were opened)
  • NDGF: Some storage network and power cables needs rerouting at NSC Tuesday 9-12 UTC. Some Atlas data unavailable. (Christian: I am having trouble connecting to Vidyo. videoportal.cern.ch is hanging.)
  • NL-T1:
    • Last week's network issue GGUS:126937 was caused by a Qfabric switch that was confused. A restart fixed it. The Qfabric platform caused many problems over the last 6 to 12 months because Juniper wants to discontinue this platform and the support leaves much to be desired.
    • SURFsara tape maintenance tonight: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=22626
    • Perhaps we won't be able to dial in.
  • NRC-KI: nc
  • OSG: mc
  • PIC: dCache upgraded last week. Minor issues need solution, 20 min downtime requested for tomorrow
  • RAL: A few low level issues.
    • The Hyper-V infrastructure at RAL failed on Thursday - We had a brief un-scheduled downtime
    • There was a squid problem over the weekend. We reset the max-files parameter and believe it is fixed.
  • TRIUMF:NTR

  • CERN computing services:
    • The two network interventions planned (15th and 22nd) have now all been scheduled for the 15th (extended in time 05.00 - 09.00) OTG0036115 at LHCb's request
    • ...and the intervention planned on 22nd has been canceled.
    • Lots of services have declared downtimes in CERN ITSSB.
  • CERN storage services:
    • FTS : Network intervention planned for 22nd rescheduled for the 15th (OTG:0036199). Expected impact: Up to three times of 7 minutes interruptions
    • Network intervention on 15th March will result in EOS, CERNBox, CASTOR, AFS services being 'degraded' (three times of 7 minutes interruptions)
    • 14th March: CASTORLHCb will be down due to DB migration to new H/W
    • 15th March: EOSCMS will be degraded due to move to new H/W during network intervention
    • 16th March: CASTOR will be down due to DB migration to new H/W
  • CERN databases:
    • 1 instance of CMSR database was rebooted on Sunday morning around 4:30 am
    • Network intervention planned for 22nd rescheduled for the 15th (OTG:0036199). Expected impact: Up to three times of 7 minutes when databases will be up but not available. All the DBs will be at risk during the intervention.
    • LCGR database will be migrated to new HW on Wednesday, 15th at 10:30, 2 hours downtime expected (OTG:0036148)
    • CMSR database will be migrated to new HW on Tuesday, 21st at 13:30, 2,5 hours downtime expected (OTG:0036368)
    • CMSONR database will be migrated to P5 in the week 27th at 09:00, 3 hours downtime expected, exact data will be announced later this week (OTG:0036369)
  • GGUS: NTR
  • Monitoring: NP
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r22 < r21 < r20 < r19 < r18 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r22 - 2017-03-14 - BelindaChan
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback